OpenStudy (anonymous):

PLEASE HELP!!!!!!!!!!!!!!!!! The sides of a rectangle are in the ratio 5:7 and the perimeter of the rectangle is 96 cm. What is the area of the rectangle? a.980 cm2 b.560 cm2 c.462 cm2 d.418 cm2

OpenStudy (anonymous):

Let length of the rectangle be l = 7x and breadth be b= 5x since Perimeter P = 2(l+b) 96 = 2(7x+5x) 96 = 2*12x 96 = 24x x = 4 therefore l = 7x = 7* 4 = 28 cm b = 5x = 5*4 = 20 cm since Area of rectangle = length * breadth = l*b= 28*20= 560 sq cm